# Keeping Up with the Models: Online Deployment and Routing of LLMs at Scale

### v1

@@ 4 StageRoute: Stage-Based LLM Deployment and Routing Algorithm
We now introduce StageRoute (Algorithm 1), a two-level hierarchy that unifies deployment and per-query routing in one algorithm: (i) Strategic layer. At each discrete update point $\tau_{k}$, the algorithm decides which models to deploy, adapting to newly available LLMs while respecting capacity and budget limits. (ii) Tactical layer. Between updates, it routes every incoming query in real time among the currently deployed models. The system starts with prior parameter estimates, an empty unexplored-model list, $\tau_{0}=0$, and an empty initial deployment $\mathcal{D}_{0}(\mathcal{A})$. It then proceeds through $K$ stages defined by update points $\{\tau_{1},\dots,\tau_{K}\}$. At the start of each state $k$ ($t=\tau_{k}$), the algorithm executes two phases in sequence: model deployment followed by request routing.
@@
Model Deployment Phase (Stage Start): At each update point $\tau_{k}$, StageRoute first adds every LLM that has become eligible since the previous update—those with $t_{m}\leq\tau_{k}$ and $t_{m}>\tau_{k-1}$—and initializes their parameter estimates. With the enlarged pool $\mathcal{M}_{\tau_{k}}$, StageRoute then solves the deployment optimization in Eq. (3) to pick the models for stage $k$.
@@
$0\leq d_{m}\leq\alpha_{m}\text{ for }m\in\mathcal{M}_{\tau_{k}},|\text{supp}(d)|=\min(M_{\max},|\mathcal{M}_{\tau_{k}}|)\Big{\}}.$
The decision vector $d^{*}=(d^{*}_{m})_{m\in\mathcal{M}_{\tau_{k}}}$ maximizes an optimistic performance surrogate while respecting budget and capacity constraints, using upper confidence bounds (UCBs) on rewards ($\mu^{U}_{m}$) and lower confidence bounds (LCBs) on costs ($c^{L}_{m}$), both derived from data up to $\tau_{k}.$
Specifically, let $\bar{\mu}_{m}(\tau_{k})$ and $\bar{c}_{m}(\tau_{k})$ be the empirical mean reward and cost of model $m$ based on $N_{m}(\tau_{k})$ selections observed up to $\tau_{k}$. Define the UCBs and LCBs as:
$c^{L}_{m}$ $:=\text{proj}_{[c_{1},c_{2}]}\left(\bar{c}_{m}(\tau_{k})-2f_{rad}(\bar{c}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)\right).$ (5)
The deployment optimization in Eq. (3) maximizes expected utility subject to the budget $b$, per-model capacity limits $\alpha_{m}$, and the cardinality cap $|\text{supp}(d)|=\min(M_{\max},|\mathcal{M}_{\tau_{k}}|)$. Its solution $d^{*}$ determines the active set for stage $k$:
The value $d^{*}_{m}$ themselves are not used as routing probabilities; they serve only to pick the most promising feasible models. The deployment set $\mathcal{D}_{k}(\mathcal{A})$ then remains fixed until the next update point $\tau_{k+1}.$
Algorithm 1 StageRoute: Stage-Based LLM Routing
0:  Update points (stage boundaries) $\{\tau_{1},\dots,\tau_{K}\}$; overall operational budget $b$; maximum number of concurrently deployed models $M_{\max}$
1:  Initialize: Prior estimates for their parameters; unexplored set $\mathcal{U}\leftarrow\emptyset$; initial time $\tau_{0}\leftarrow 0$; initial deployed set $\mathcal{D}_{0}(\mathcal{A})\leftarrow\emptyset$
2:  ///// /Stage-wise Deployment Phase:
3:  for each stage $k=1$ to $K$ do
4:     Add newly available models $\{m\mid t_{m}\leq\tau_{k}\text{ and }t_{m}>\tau_{k-1}\}$, and initialize their parameters
5:     Determine the deployment decision vector $d^{*}=(d^{*}_{m})_{m\in\mathcal{M}_{\tau_{k}}}$ by solving DeployOPT (3)
6:     Update the set of deployed models for stage $k$: $\mathcal{D}_{k}(\mathcal{A})\leftarrow\{m\in\mathcal{M}_{\tau_{k}}\mid d^{*}_{m}>0\}$
7:     ///// /Per-query Routing Phase (for query at time $t$):
8:     for each time step $t$ from $\tau_{k}$ to $\tau_{k+1}-1$ (if $k<K$), or to end of horizon (if $k=K$do
9:        Compute the optimal routing distribution $p_{t}^{*}=(p_{t}^{*}(m))_{m\in\mathcal{D}_{k}(\mathcal{A})}$ by solving RouteOPT (7)
10:        Sample model $m_{t}\sim p_{t}^{*}$, and route request to the chosen model $m_{t}$
11:        Observe the actual reward $r_{t}$ and actual cost $c_{t}$ resulting from using model $m_{t}$
12:        Update the parameters for model $m_{t}$
13:     end for
14:  end for
Request Routing Phase (Intra-Stage): After each deployment update, StageRoute handles each query at time $t\in[\tau_{k},\tau_{k+1}-1)$ (or until $T$ if $k=K$) in four steps as follows: (1) Routing LP. It solves the linear program (LP) in Eq. (7) to compute the optimal routing distribution $p_{t}^{*}=(p_{t}^{*}(m))_{m\in\mathcal{D}_{k}(\mathcal{A})}$ over the currently deployed models $\mathcal{D}_{k}(\mathcal{A})$.
[…]
$\quad\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}p_{t}(m)=1,0\leq p_{t}(m)\leq\alpha_{m}\text{ for }m\in\mathcal{D}_{k}(\mathcal{A})\Big{\}}.$
@@
This LP maximizes the expected reward by combining the current UCBs for reward ($\mu^{U}_{m}$) and LCBs for cost ($c^{L}_{m}$) while enforcing the per-query budget $b$ (through the cost bounds) and each model’s capacity limit $\alpha_{m};$ the cardinality cap $M_{\max}$ is already respected because $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$. (2) Model selection. A model $m_{t}$ is sampled according to $p_{t}^{*}$ and used to serve the query. (3) Feedback. StageRoute observes the realized reward $r_{t}$ and cost $c_{t}$. (4) Update statistics. StageRoute then updates $\mu_{m_{t}}$, $c_{m_{t}}$ and $N_{m_{t}}$, recomputing the corresponding confidence bounds so that future routing—and the next deployment decision—reflect the latest information.
Following the deployment decision, the algorithm enters an inner loop that processes each query arriving at time $t$ from $\tau_{k}$ to $\tau_{k+1}-1$ (or to the end of the horizon if $k=K$). For each query, the algorithm performs the following steps: It determines the optimal routing probability distribution $p_{t}^{*}=(p_{t}^{*}(m))_{m\in\mathcal{D}_{k}(\mathcal{A})}$ over the currently deployed models $\mathcal{D}_{k}(\mathcal{A})$ by solving the linear program (LP) in Equation (7). This LP maximizes the expected reward based on current UCBs for reward ($\mu^{U}_{m}$) and LCBs for cost ($c^{L}_{m}$) for models in $\mathcal{D}_{k}(\mathcal{A})$. These UCB/LCB estimates are based on all information gathered up to the previous time step $t-1$ and are updated as new observations are made. The routing LP adheres to the long-term budget $b$ (applied on a per-query basis using LCBs for costs) and individual model capacity constraints $\alpha_{m}$ for $m\in\mathcal{D}_{k}(\mathcal{A})$. Note that the $M_{\max}$ constraint is not part of this LP, as the set $\mathcal{D}_{k}(\mathcal{A})$ already satisfies $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$ (due to the constraint in Equation 3 and the formation of $\mathcal{D}_{k}(\mathcal{A})$). A model $m_{t}$ is then sampled according to this distribution $p_{t}^{*}$ (Line 10), and the incoming query is routed to $m_{t}$. The system observes the actual performance (reward) $r_{t}$ and cost $c_{t}$ from the selected model $m_{t}$. Finally, this feedback is used to update the parameters for model $m_{t}$, such as its empirical average reward, empirical average cost, and usage count. These updated statistics are then used to refine the $\mu^{U}_{m_{t}}$ and $c^{L}_{m_{t}}$ estimates for subsequent routing decisions and for the deployment decision at the next stage.
5 Theoretical Analysis
@@
We analyze StageRoute (Algorithm 1) by deriving an upper bound on its cumulative regret and a matching lower bound that applies to any online algorithm. Taken together, these results show that StageRoute is near-optimal in the worst case.
@@ 5.1 Upper Bound
Consider StageRoute running for $T$ queries divided into $K$ stages, with $N=|\mathcal{M}_{T}|$ models that may arrive over time. Set the confidence parameter to $\gamma=\Theta(\log(NT/\delta))$ to obtain overall confidence $1-\delta$. Then StageRoute’s expected regret satisfies
@@
$\textnormal{\text{Regret}}(\textnormal{{StageRoute}})\leq\mathcal{O}\left(\sqrt{M_{\max}KT\log(NT/\delta)}+\frac{NT}{M_{\max}K}\right).$
@@
When $K=\Omega(T^{1/3})$ and $M_{\max}=\Omega(N^{2/3})$, we have $\textnormal{\text{Regret}}(\textnormal{{StageRoute}})\leq\tilde{\mathcal{O}}\left(N^{1/3}T^{2/3}\right).$
The theorem’s primary regret bound demonstrates that overall system performance is governed by a trade-off between two key components: The regret bound reveals a tradeoff between two components. The statistical term $\mathcal{O}(\sqrt{M_{\max}KT\log(NT/\delta)})$ reflects the difficulty of learning optimal routing and deployment. The structural term $\mathcal{O}(NT/(M_{\max}K))$ captures the exploration bottleneck created by probing a large universe of $N$ models through only $M_{\max}$ deployment slots over $K$ stages. This tension means that both the number of stages $K$ and the deployment capacity $M_{\max}$ must be tuned carefully. When chosen well, the overall regret matches the lower bound in Theorem  2, yielding a near-optimal order of $\widetilde{\mathcal{O}}(T^{2/3})$. Although prior work suggests that good routing can rely on a small set of models, our analysis shows this is not true in dynamic environments with continual model turnover. In such settings, maintaining a sufficiently large concurrent deployment capacity is crucial for near-optimal performance.
Our analysis unifies stage-wise decision making with a continually evolving model pool, tackling the practical need to deploy newly released LLMs under tight capacity limits. We expose a new model-discovery bottleneck that captures the exploration challenge created by the mismatch between the limited number of concurrent slots ($M_{\max}$) and the discrete schedule of deployment stages. This bottleneck is distinct from the usual learning-related regret incurred by routing and deployment decisions. The theory shows that near-optimal performance requires setting the stage count to $K=\Theta(T^{1/3})$ and provisioning at least $M_{\max}=\Omega(N^{2/3})$ concurrent slots. These prescriptions offer actionable guidance for system designers and can inform future algorithmic and infrastructure choices in real-world LLM-serving environments.
Proof Sketch..
We decompose the algorithm’s total expected regret into two parts, i.e. the expected routing regret, $\mathcal{R}_{\text{routing}}(T)$ and the expected deployment regret, $\mathcal{R}_{\text{deploy}}(T)$, so that
$\text{Regret}(\mathcal{A})=\underbrace{\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{M}_{\tau_{k}})-V(b,\mathcal{D}_{k})\right]}_{\mathcal{R}_{\text{deploy}}(T)}+\underbrace{\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{D}_{k})-r_{t}\right]}_{\mathcal{R}_{\text{routing}}(T)}.$
The routing regret, $\mathcal{R}_{\text{routing}}(T)$, arises because the algorithm relies on estimated model parameters when routing queries within each deployed set $\mathcal{D}_{k}$. For a single stage $k$, the expected routing regret is bounded by $\mathcal{O}\left(\sqrt{{\gamma}|\mathcal{D}_{k}|T\cdot V(b,\mathcal{D}_{k})/K}\right)$. Summing the regret bounds over all $K$ stages, the total expected routing regret is bounded by
$\mathcal{R}_{\text{routing}}(T)\leq\mathcal{O}\left(\sqrt{\gamma M_{\max}KT}\right).$
The deployment regret, $\mathcal{R}_{\text{deploy}}(T)$, is more intricate and has two component. (i) Learning error: uncertainty in the model estimates when selecting the deployment set $\mathcal{D}_{k}$. (ii) Model discovery bottleneck: the $M_{\max}$-slot limit can delay evaluation and adoption of newly arrived, high-quality models. The sum of these two terms is bounded by
$\mathcal{R}_{\text{deploy}}(T)\leq\mathcal{O}\left(\sqrt{\gamma T\cdot\min(N,KM_{\max})}+\frac{NT}{M_{\max}K}\right).$
Combining the routing- and deployment-regret bounds and substituting $\gamma=\Theta(\log(NT/\delta))$ gives the overall regret bound stated in the theorem; the full proof appears in Appendix B. ∎

### v2

@@ 3 StageRoute: Stage-Based LLM Deployment and Routing
We introduce StageRoute (Algorithm 1), a two-level hierarchy that unifies deployment and per-query routing in one algorithm: (i) Strategic layer. At each discrete update point $\tau_{k}$, the algorithm decides which models to deploy, adapting to newly available LLMs while respecting the budget and operational constraints. (ii) Tactical layer. Between updates, it routes every incoming query in real time among the currently deployed models. The system starts with prior parameter estimates, an empty unexplored-model list, $\tau_{0}=0$, and an empty initial deployment $\mathcal{D}_{0}(\mathcal{A})$. It then proceeds through $K$ stages, and at the start of each stage $k$ ($t=\tau_{k}$), the algorithm executes two phases in sequence: model deployment followed by request routing.
Algorithm 1 StageRoute: stage-based LLM deployment (active-set selection) and online query routing
0: Update points $\{\tau_{1},\dots,\tau_{K}\}$; budget $b$; concurrency cap $M_{\max}$
1:Initialize: Prior parameter estimates; $\tau_{0}\leftarrow 0$; $\mathcal{D}_{0}(\mathcal{A})\leftarrow\emptyset$
2:////Stage-wise Deployment Phase:
3:for $k=1$ to $K$ do
4:  Incorporate newly available models $\{m\mid t_{m}\leq\tau_{k}<t_{m}\!+\!(\tau_{k}\!-\!\tau_{k-1})\}$; initialize their parameters
5:  Solve DeployOPT (3) for $d^{*}$ and set $\mathcal{D}_{k}(\mathcal{A})\leftarrow\{m\mid d^{*}_{m}>0\}$
6:  ////Per-query Routing Phase (for query at time $t$):
7:  for $t=\tau_{k}$ to $\tau_{k+1}-1$ (or to $T$ if $k=K$) do
8:   Compute routing distribution $p_{t}^{*}$ by solving RouteOPT (7)
9:   Sample $m_{t}\sim p_{t}^{*}$ and route the query to it
10:   Observe reward $r_{t}$ and cost $c_{t}$; update statistics for $m_{t}$
11:  end for
12:end for
@@
Model Deployment Phase (Stage Start). At each update point $\tau_{k}$, StageRoute first incorporates any newly available models (those with $\tau_{k-1}<t_{m}\leq\tau_{k}$) and initializes their parameter estimates. With the enlarged pool $\mathcal{M}_{\tau_{k}}$, StageRoute solves the deployment optimization in Eq. (3) to pick the models for stage $k$:
DeployOPT: $\max_{d\in\Delta(\mathcal{M}_{\tau_{k}})}\Big\{\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu^{U}_{m}d_{m}\;\Big|\;\sum_{m\in\mathcal{M}_{\tau_{k}}}c^{L}_{m}d_{m}\leq b,\sum_{m\in\mathcal{M}_{\tau_{k}}}d_{m}=1,$ (3)
$\qquad\quad 0\leq d_{m}\leq\alpha_{m}\text{ for }m\in\mathcal{M}_{\tau_{k}},|\text{supp}(d)|=\min(M_{\max},|\mathcal{M}_{\tau_{k}}|)\Big\}.$
This optimization problem is a Mixed-Integer Program (MIP). The combinatorial nature arises from the cardinality constraint on the support of $d$, which limits the number of active models. A standard way to formulate this is by introducing a binary activation variable $z_{m}\in\{0,1\}$ for each model. The full stage-$k$ deployment problem can then be written as:
$\max_{d,z}\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu^{U}_{m}d_{m},\quad\text{s.t.}\quad$ $\sum_{m}c^{L}_{m}d_{m}\leq b,~\sum_{m}d_{m}=1,~0\leq d_{m}\leq\alpha_{m}z_{m},$
$~\sum_{m}z_{m}=\min(M_{\max},|\mathcal{M}_{\tau_{k}}|),~z_{m}\in\{0,1\}.$
Here, the binary variables $z_{m}$ explicitly select which models are live for the stage, while the continuous variables $d_{m}$ represent an optimistic deployment mix. The solution to this MIP $d^{*}$ maximizes an optimistic performance surrogate using UCBs for rewards ($\mu^{U}_{m}$) and LCBs for costs ($c^{L}_{m}$), which are derived from data up to $\tau_{k}$. Specifically, let $\bar{\mu}_{m}(\tau_{k})$ and $\bar{c}_{m}(\tau_{k})$ be the empirical mean reward and cost of model $m$ based on $N_{m}(\tau_{k})$ selections observed up to $\tau_{k}$. Define the UCBs and LCBs as:
$c^{L}_{m}$ $:=\text{proj}_{[c_{1},c_{2}]}\left(\bar{c}_{m}(\tau_{k})-2f_{rad}(\bar{c}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)\right),$ (5)
The deployment optimization in Eq. (3) maximizes expected utility subject to the budget $b$, per-model throughput limits $\alpha_{m}$, and the concurrency cap $|\text{supp}(d)|=\min(M_{\max},|\mathcal{M}_{\tau_{k}}|)$. Its solution $d^{*}$ determines the active set for stage $k$:
Crucially, the values $d^{*}_{m}$ are not used as routing probabilities; they serve only to select the most promising feasible models. The deployment set $\mathcal{D}_{k}(\mathcal{A})$ remains fixed until the next update point $\tau_{k+1}$.
Request Routing Phase (Intra-Stage). For each query arriving at time $t\in[\tau_{k},\tau_{k+1})$, StageRoute performs four steps: (1) Routing LP. It solves the linear program (LP) in Eq. (7) to compute the optimal routing distribution $p_{t}^{*}=(p_{t}^{*}(m))_{m\in\mathcal{D}_{k}(\mathcal{A})}$ over the currently deployed models $\mathcal{D}_{k}(\mathcal{A})$.
$\texttt{RouteOPT:}\quad\max_{p_{t}\in\Delta(\mathcal{D}_{k}(\mathcal{A}))}\Big\{$ $\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\mu^{U}_{m}p_{t}(m)\;\Big|\;\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}c^{L}_{m}p_{t}(m)\leq b,$ (7)
@@
This LP maximizes the expected reward by combining the current UCBs for reward ($\mu^{U}_{m}$) and LCBs for cost ($c^{L}_{m}$) while enforcing the per-query budget $b$ (through the cost bounds) and each model’s throughput limit $\alpha_{m}$. (2) Model selection. Sample $m_{t}\sim p_{t}^{*}$ and serve the query. (3) Feedback. StageRoute observes the realized reward $r_{t}$ and cost $c_{t}$. (4) Update. Refresh $\bar{\mu}_{m_{t}}$, $\bar{c}_{m_{t}}$, $N_{m_{t}}$ and recompute $\mu^{U}_{m_{t}},c^{L}_{m_{t}}$ for subsequent routing and the next deployment decision.
Algorithmic Innovations. While StageRoute builds on the principle of “optimism in the face of uncertainty”, its architecture is tailored to dynamic LLM deployment and departs from standard bandit formulations. First, it imposes a hierarchical decision structure that mirrors operational practice: the deployment decision is a high-stakes, combinatorial choice whose consequences persist for an entire stage, a form of long-term commitment absent from standard, per-round bandits. Second, staged updates create a structured delay in acting on feedback. Information about non-deployed models cannot influence decisions until the next stage, inducing an exploration-exploitation tradeoff that requires anticipating performance over the whole stage, not just the next round. Finally, StageRoute decouples deployment from routing execution: DeployOPT determines only the active set $\mathcal{D}_{k}(\mathcal{A})$, while the per-query policy is recomputed online via RouteOPT. This separation enables rapid query-level adaptation even when the underlying infrastructure remains fixed during a stage.
4 Theoretical Results
@@
We analyze StageRoute (Algorithm 1) by deriving an upper bound on its cumulative regret and a matching lower bound that applies to any online algorithm for this problem. Together, these results show that StageRoute is near-optimal in the worst case.
@@ 4.1 Upper Bound
Consider StageRoute running for $T$ queries divided into $K$ stages, with a concurrency cap $M_{\max}$ and $N=|\mathcal{M}_{T}|$ total models arriving over time. Set the confidence parameter to $\gamma=\Theta(\log(NT/\delta))$ to obtain overall confidence $1-\delta$. Then the expected regret is bounded by:
@@
$\textnormal{Regret}(\textnormal{{StageRoute}})\leq\mathcal{O}\left(\sqrt{M_{\max}KT\log(NT/\delta)}+\frac{NT}{M_{\max}K}\right).$
@@
Choosing $K=\Theta(T^{1/3})$ and $M_{\max}=\Omega(N^{2/3})$ yields $\textnormal{Regret}(\textnormal{{StageRoute}})\leq\widetilde{\mathcal{O}}\left(N^{1/3}T^{2/3}\right)$.
The two terms in the bound reflect complementary sources of difficulty. The first term, $\widetilde{\mathcal{O}}\big(\sqrt{M_{\max}KT}\big)$, is the statistical learning cost of routing within the deployed set. The second term, $\widetilde{\mathcal{O}}\big(NT/(M_{\max}K)\big)$, is a structural model-discovery bottleneck: when only $M_{\max}$ models can be active at a time across $K$ stages, exploration is throttled. Strong late-arriving models can be missed unless sufficient deployment slots and update frequency are provisioned. Balancing these two terms gives the near-optimal $\widetilde{\mathcal{O}}(T^{2/3})$ rate, which matches the lower bound in Theorem 2.
Practical guidance. The bound yields an actionable rule: to approach the optimal rate, set the number of stages to approximately $K\approx T^{1/3}$ (implying a stage length of $T^{2/3}$) and provision a concurrency cap $M_{\max}$ large enough to track new arrivals (ideally $M_{\max}\approx N^{2/3}$ when feasible). However, updating more frequently ($K\gg T^{1/3}$) is also ill-advised. Even if it does not increase regret, it cannot offer further asymptotic improvement due to the lower bound, while needlessly incurring significant computational and operational overhead with each additional deployment stage. This underscores that exploration capacity, defined by the concurrency cap and update frequency, is itself a scarce system resource to be optimized. Relying on only a few “top” models is provably suboptimal in a dynamic model pool.
Why existing analyses do not apply. Classical MAB and BwK typically assume per-round choices from a static set and lack a hard concurrency cap; CMAB selects superarms from a fixed base set and does not model stage-level commitment; streaming bandits allow arrivals but do not couple stage-committed deployment with budget and per-model throughput constraints. Our setting is distinct due to: (i) a concurrency cap that explicitly constrains the support of the deployment optimization, (ii) stage-committed deployment that induces a structured delay in acting on new information, and (iii) the simultaneous enforcement of a long-term budget and per-model throughput limits.
Proof ideas and new technical elements. Our proof introduces a virtual optimal deployment set to bridge the offline benchmark and the online policy, yielding a clean regret decomposition into stage-level deployment regret and intra-stage routing regret. The routing term is handled with standard confidence arguments. The deployment term requires two new ingredients: (a) a quantification of the model-discovery bottleneck caused by the limited concurrency cap and discrete updates (showing how the $NT/(M_{\max}K)$ term arises), and (b) a support-aware sensitivity analysis of the deployment LP (via its dual), bounding how UCB/LCB estimation errors perturb the optimal active set under the concurrency constraint. Together, these yield Theorem 1. These elements differ fundamentally from standard learning-regret analyses and may inform further work on staged, combinatorial online decision problems. Complete details appear in Appendix C.